In six years time I will be one and a quarter times as old as i was four years ago, How old is he?
Answer by lwsshak3(11628)

In six years time I will be one and a quarter times as old as i was four years ago, How old is he?
***
x=present age
..
in 6 years:
x+6
..
4 yrs ago
x-4
..
x+6=5/4(x-4)
4x+24=5x-20
x=44
How old is he? 44